OpenStudy (anonymous):

Determine a. if f has a maximum value or a minimum value; b. the value of x at which the maximum or minimum occurs; c. the maximum or minimum value of f: f(x)=2x^2+3x+2

myininaya (myininaya):

since the coefficient of x^2 is postive then the parabola opens upward so U that means we have a minmum

OpenStudy (anonymous):

Oh, wow, now that made sense...so the first part is that easy..okay, i'm following thoroughly

myininaya (myininaya):

f(x)=2(x^2+3x/2)+2 f(x)=2(x^2+3x/2+(3/4)^2)+2-2(3/4)^2 f(x)=2(x+3/4)^2+2-9/8 f(x)=2(x+3/4)^2+7/8 so the max is x=-3/4 the max value is 7/8

OpenStudy (anonymous):

wow...that made sense...not too bad actually

myininaya (myininaya):

lol usually people get confused when I add in a weird zero

OpenStudy (anonymous):

I didn't see a weird 0

OpenStudy (anonymous):

lol..thanks again

myininaya (myininaya):

ur wecolme